### v1

@@ Abstract
The era of exascale computing presents both exciting opportunities and unique challenges for quantum mechanical simulations. While the transition from petaflops to exascale computing has been marked by a steady increase in computational power, the shift towards heterogeneous architectures, particularly the dominant role of graphical processing units (GPUs), demands a fundamental shift in software development strategies. This review examines the changing landscape of hardware and software for exascale computing, highlighting the limitations of traditional algorithms and software implementations in light of the increasing use of heterogeneous architectures in high-end systems. We discuss the challenges of adapting quantum chemistry software to these new architectures, including the fragmentation of the software stack, the need for more efficient algorithms (including reduced precision versions) tailored for GPUs, and the importance of developing standardized libraries and programming models.
keywords:
Quantum Mechanical Simulations, High-Performance Computing, Heterogeneous Architectures, Software Stack, Exascale1 Rationale and scope
Until recently, the trajectory of top-performing computers has been fairly predictable, guided by a handful of leading hardware vendors [1]. These vendors have consistently provided scientists with clear paths (via stable languages, compilers, libraries, and tools) to harness escalating computational power, culminating in the era of petaflops [2, 3, 4]. The move to exascale computing comes with unique challenges, including a rapidly changing array of hardware options [5, 6, 7, 8, 9] and the once-reliable software stack that is now becoming increasingly fragmented and inconsistent [10, 11, 12, 13, 14, 15, 16]. After traditional applications like cryptocurrency mining and gaming, artificial intelligence (AI) has effectively harnessed the power of Graphical Processing Units (GPUs). This success is due to a strong alignment between AI algorithms and the architectural strengths of GPUs, coupled with the development of specialized libraries [17, 18, 19, 20]. The transition to exascale for quantum mechanical simulations uncovers an intricate web of prospects and predicaments for physicists and chemists [21, 4, 22] and typically requires large community efforts (e.g. US Exascale Computing Project [23], EU Centers of Excellence [24] and National Competence Centers [25]) well beyond individual capabilities. This review explores the uneven path toward exascale simulations, highlighting both promising developments and persistent challenges [26, 27, 28, 29, 30, 31, 32, 33] and offering a nuanced understanding of this important shift in computational science.
2 The transition to exascale
@@
Quantum mechanical (QM) simulations are essential for understanding and predicting the behavior of complex materials and molecular systems across various scientific disciplines. The accuracy of these simulations hinges on precisely describing electronic structures, a task that is computationally demanding and has been dramatically enhanced by advancements in high-performance computing (HPC). The progress in HPC is marked by milestones tracked by the Top500 list [34], which monitors peak computational power through a benchmark known as LINPACK [35]. This algorithm measures the performance of systems in solving a dense linear system, a task optimized for high computational throughput with minimal data movement and communication. Remarkably, LINPACK achieves around 70% of the theoretical maximum speed, derived from the total number of cores multiplied by their nominal speed. Figure 1(1(a)) exhibits the steady progress of computational power, culminating in crossing the exaflop barrier approximately two years ago.
@@
However, LINPACK’s numbers, while impressive, are not fully representative of real-world performance. A more challenging and realistic benchmark is the High-Performance Conjugate-Gradient (HPCG) test, which evaluates the ability to solve sparse linear systems and is therefore representative of a different type of numerical applications, for instance, based on finite elements and finite difference schemes [36]. The performance here is starkly different, typically about 20 times lower than LINPACK, utilizing less than 5% of the nominal speed [37]. This contrast underscores the difference between idealized benchmarks and practical workloads.
@@
While LINPACK and HPCG represent reasonable upper and lower bounds of performance, respectively, neither reflects the complexity of actual applications. The Gordon Bell Prize, awarded for outstanding achievements in high-performance computing applications, provides a more relevant measure [38]. These applications, often solving real scientific problems, have demonstrated performance nearing the exaflop threshold [39, 40]. Achieving this requires monumental efforts: entire codebases are rewritten by interdisciplinary teams of physicists, chemists, mathematicians, and computer scientists. New algorithms are developed, and implementations are tailored to exploit the specific features of the target architectures. In contrast, legacy codes, which constitute the bulk of real-world applications, lag significantly behind. These codes were not designed with exascale capabilities in mind and thus fall short of the exaflop performance. The challenge ahead lies in bridging this gap, ensuring that applications needing exascale capabilities can harness the full potential of exascale systems.
3 Hardware: The changing landscape of computational power
@@
The transition from petaflops to exascale computing marked a significant departure from the consistent and predictable trajectory of high-performance computing. While the climb to petaflops was characterized by steady increases in processor core count and clock speeds, exascale computing brings a fundamental shift in hardware architectures and software development strategies. One of the most significant changes is the increasing adoption of heterogeneous architectures. Unlike the homogeneous central processing unit (CPU)-based systems that dominated petaflops, exascale systems heavily rely on accelerators like GPUs, which offer massive parallelism for specific tasks but come with unique challenges. This heterogeneous landscape presents a major hurdle for software developers as traditional algorithms and software stacks designed for CPUs need significant modifications to utilize the capabilities of GPUs efficiently.
@@
Before exploring this hardware evolution, it is crucial to identify the primary application areas driving these advancements, as they fundamentally determine the level of investment and the trajectory of hardware development. Scientific HPC, including QM applications, remains a niche market and does not significantly influence major hardware innovations. Instead, scientific computing and exaflop architectures often repurpose software and hardware technology developed for other, larger markets. Two key segments, that are increasingly shaping the practices of the scientific community, are cloud and accelerator computing, driven primarily by their adoption in conventional companies and the field of AI, respectively. The general-purpose laptop segment also indirectly impacts high-end processor computing. For instance, Apple, a major customer of the Taiwan Semiconductor Manufacturing Company (TSMC), exerts significant pressure on high-end chip production for both cloud computing and AI applications due to their high volume demand, as these components share the same fabrication lines [41].
@@
On the CPU front, changes have been incremental over the past decade. Major CPU manufacturers (AMD, Intel, ARM) have converged towards a similar generic architecture: multicore processors based on out-of-order and superscalar technology. A significant development has been the increase in the number of cores per processor, now exceeding one hundred. While other technologies have emerged, none have achieved widespread adoption. Wider vectors (up to 512 bits) have been embraced by x86 architectures (initially by Intel, followed by AMD), whereas ARM continues to use shorter vectors (128/256 bits) without major performance penalties. Notably, memory technology has evolved more significantly, with very large level 3 caches using three-dimensional stacking and High Bandwidth Memory (HBM) offering an alternative to the standard Double Data Rate Synchronous Dynamic Random Access Memory (DDR-SDRAM). The performance gains from these new memory technologies depend heavily on the application’s data access characteristics.
Nvidia H100 SXM | AMD EPYC™ 9754 | ||
16896 cores, 1.6 GHz | 128 cores, 2.25 GHz | ||
Release date | September 2022 | June 2023 | |
FP64 | 34 TFlop/s | 6.9 TFlop/s | |
FP64 (tensor core) | 67 TFlop/s | ||
FP32 | 67 TFlop/s | 13.8 TFlop/s | |
FP32 (tensor core) | 989 TFlop/s | ||
FP16 (tensor core) | 1979 TFlop/s | ||
FP8 (tensor core) | 3958 TFlop/s | ||
Memory | 80 GB | max 6000 GB | |
Memory Bandwidth | 3350 GB/s (HBM3) | 460.8 GB/s (DDR5) | |
Interconnection Nvlink | 900 GB/s | ||
Interconnection PCIe | 128 GB/s | ||
Quantum X800 Infiniband network | 100 GB/s | 100 GB/s | |
@@
On the accelerator front, the primary advancement has been the massive increase in core count, now reaching thousands – nearly two orders of magnitude more than CPUs. The number of cores and peak performance depend on the floating-point (FP) format; typically, double-precision (DP) cores are half as numerous as single-precision (SP) cores. The shift toward narrower FP formats has been driven by AI, which tolerates reduced precision (16-bit, 8-bit, and even 4-bit formats are in use). Modern GPUs achieve remarkable performance using smaller FP formats and Tensor Cores, which are optimized for dense matrix multiplications. To illustrate these differences, Table 1 compares the computational performance of a high-performance CPU and a GPU released within the same time frame. Peak performance reflects execution speed only when computation is the bottleneck, not memory transfers or inter-node communications, which are too slow to keep up with the execution units. With standard DP operations (FP64), GPU peak performance is about five times that of a high-end CPU. When Tensor Cores are used, this increases to a factor of ten. For SP (FP32) operations, both CPU and GPU performance doubles, but with Tensor Cores, the GPU reaches 989 TFlop/s, over 70 times that of the CPU. Reducing precision to 8 bits (FP8) adds another factor of four. While smaller FP formats can sometimes be used for standard FP32 and FP64 tasks, Tensor Cores mainly benefit computations reliant on dense matrix multiplication. Achieving GPU peak performance is harder than on CPUs, as it requires high levels of parallelism. Consequently, matrices on the GPU must be much larger [44], or numerous, requiring batching algorithms. Furthermore, GPUs have less memory than CPUs [45] and large workloads necessitate memory transfers from the host to the accelerator, which must be managed in parallel with computations to prevent communication bottlenecks. We further elaborate below on the implications of these CPU and GPU performance differences for quantum simulations.
@@
Finally, on the supercomputing systems front, the changes have been fairly radical, using Top500 as a reference, most of the major changes can be easily monitored. First, the amount of parallelism has drastically increased: the top runner in Top500 had 200 000 cores in June 2010, and now it has 9 million cores (a 45 increase). The overall system organization has also radically changed: in 2010, GPUs were used in less than 40 systems among the 500, now over half of the systems are equipped with GPUs. Today’s top systems exhibit a heterogeneous architecture, integrating both CPUs and GPUs, though the bulk of computational capability predominantly stems from the GPUs. Moreover, the total power consumption of these systems has increased by an order of magnitude, rising from 3 MW to 30 MW, albeit with a notably irregular progression, as depicted in Figure 1(1(b)). This underscores the substantial enhancements in system design, particularly in terms of energy efficiency.
4 Software: The challenge of increasing fragmentation
@@
OpenCL was introduced to address portability, providing an open standard for CPUs, GPUs, and field programmable gate arrays (FPGAs) across vendors. However, it has not gained wide adoption in HPC due to performance and usability issues. While OpenCL offers portability, its abstraction often results in suboptimal performance compared to specialized frameworks like CUDA. Its lower-level programming model is also cumbersome, requiring significant manual effort for memory management and kernel optimization. Fragmented implementations across vendors further undermine OpenCL’s portability, as performance and features can vary, forcing developers to fine-tune code for each platform.
The lack of an efficient standardized programming model for GPUs has resulted in a number of issues for developers. Firstly, it creates vendor lock-in, where codes written for one platform cannot easily be migrated to another without substantial code refactoring. For example, adopting NVIDIA programming models involves significant trade-offs between performance and portability across different hardware environments [46]. Another issue stemming from this fragmentation is the lack of compiler support for all programming models on all platforms. For instance, (Fig. 2), while CUDA is well supported on Nvidia GPUs through the CUDA Toolkit, its support on AMD and Intel hardware is limited and relies on indirect means like translation layers or third-party libraries and utilities [47, 48]. This forces developers to rely on vendor-specific compilers, which often lack the universality and feature-richness of their CPU counterparts.
@@
The lack of standardization in GPU programming has also led to a variety of community-driven, higher-level models that aim to abstract away vendor-specific details and provide a more portable programming experience. Examples of such models include Kokkos [50], RAJA [51], Alpaka [52], and, to a lesser extent, hipSYCL [53] which is limited to AMD and Intel. These abstraction models often utilize vendor-native infrastructure in the background, enabling developers to write code that can be deployed on multiple platforms without significant code changes. However, the support and standardization of these higher-level models can vary significantly, and relying on community-driven efforts for critical software infrastructures can introduce uncertainties and complexities in the long run. Comparing the performance of various GPU programming models on the LUMI supercomputer [54] reveals the potential of community-driven models like Kokkos in achieving portable performance but also highlights the need for further development and community support to ensure their long-term viability.
@@
Another important aspect of the software stack is the role of low-level libraries for key computational kernels, such as linear algebra, fast Fourier transforms, and communication routines. These libraries play a critical role in achieving high performance and efficiency on exascale systems. However, the use of different libraries across different platforms can also contribute to software fragmentation. For example, while Nvidia provides the highly optimized cuBLAS library for linear algebra operations on their GPUs, AMD offers the RocBLAS library. These libraries, despite providing similar functionality, often have different APIs and performance characteristics, further hindering code portability.

### v2

@@ Abstract
The era of exascale computing presents both exciting opportunities and unique challenges for quantum mechanical simulations. Although the transition from petaflops to exascale computing has been marked by a steady increase in computational power, it is accompanied by a shift towards heterogeneous architectures, with graphical processing units (GPUs) in particular gaining a dominant role. The exascale era therefore demands a fundamental shift in software development strategies. This review examines the changing landscape of hardware and software for exascale computing, highlighting the limitations of traditional algorithms and software implementations in light of the increasing use of heterogeneous architectures in high-end systems. We discuss the challenges of adapting quantum chemistry software to these new architectures, including the fragmentation of the software stack, the need for more efficient algorithms (including reduced precision versions) tailored for GPUs, and the importance of developing standardized libraries and programming models.
Table-of-contents summary
[H1] Introduction
Quantum mechanical (QM) simulations are crucial for understanding and predicting the behaviour of molecules and materials, offering insights that drive advancements in diverse fields from quantum chemistry to materials science. These simulations, which rely on solving the electronic Schrödinger equation, are computationally demanding, and their accuracy is directly tied to the complexity of the system being modeled and the chosen level of theory. Efficient execution of QM simulations has traditionally been enabled by advances in high-performance computing (HPC), with a relatively predictable trajectory of hardware and software development. A few dominant hardware vendors provided stable tools and clear paths for scientists to take advantage of increasing computational power, primarily through more central processing unit (CPU) cores and higher clock speeds. This steady progress culminated in the petascale era [1, 2, 3, 4].
However, the shift to exascale computing has disrupted this trajectory, representing not merely a quantitative leap with a thousand-fold increase in computational power, but a qualitative shift in the landscape of HPC [5, 6, 7, 8, 9]. Instead of simply scaling existing architectures, the exascale era is defined by heterogeneous architectures. These architectures prominently feature graphical processing units (GPUs) alongside traditional CPUs. This mirrors the successful utilization of GPU power by artificial intelligence (AI) applications, thanks to the alignment of AI algorithms with GPU strengths and the development of specialized libraries [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. This shift presents both exciting prospects and significant challenges for physicists and chemists performing QM simulations [21, 3, 22].
A major issue is the increasing fragmentation of the once reliable software stack, making it difficult to efficiently utilize these new, diverse hardware options [14, 15, 16, 17, 18, 19, 20]. Existing QM codes and algorithms, often optimized for CPU-based systems, are not readily transferable to GPU-accelerated environments. Addressing this requires substantial, coordinated efforts, often exceeding individual capabilities and requiring large community initiatives such as the US Exascale Computing Project [23], EU Centers of Excellence [24], and National Competence Centers [25]. The development and adoption of modular standardized libraries for key computational kernels will be crucial to simplify code development and ensure long-term sustainability [26, 27].
This review explores the uneven path toward exascale QM simulations, examining the implications of the changing hardware landscape, particularly the rise of heterogeneous architectures. We will discuss how fragmentation of the software stack impacts the development and deployment of QM software, and highlight the need for new algorithms and programming models that can effectively exploit the capabilities of GPUs, including the use of reduced precision calculations. Concrete examples from a wide range of QM methodologies will illustrate these issues.
Finally, we argue that a successful transition requires a collaborative and multifaceted approach. This includes designing new algorithms optimized for heterogeneous architectures, promoting energy-efficient computing practices, and co-designing codes with stronger collaboration between academia and industry. By addressing these challenges, the field can fully leverage exascale computing [28, 29, 30, 31, 32, 33, 34, 35], and unlock new frontiers in materials design, molecular discovery, and a deeper understanding of the quantum world.
[H1] Quantum mechanical simulations
The computational cost of QM simulations is closely tied to the number of basis functions used to represent the wave function, which scales with the size and complexity of the system. Additionally, the choice of method for solving the electronic Schrödinger equation determines both the computational scaling and memory requirements. However, despite algorithmic and HPC advancements, QM simulations on the whole remain computationally demanding, even with the advent of exascale computing.
Mean-field-like methods, including Hartree–Fock (HF), Kohn–Sham density functional theory (Kohn–Sham DFT) [36], and density functional tight-binding (DFTB) [37], rely on self-consistent iterations to optimize the density matrix and minimize the total energy. In each iteration, the Fock or Kohn–Sham matrix is constructed using one- and two-electron integrals, followed by matrix diagonalization to update molecular orbitals. These steps, particularly the construction and diagonalization of the Fock or Kohn–Sham matrix, represent the main computational bottlenecks. The matrix diagonalization scales as , and the computation of two-electron integrals scales as , making direct storage (for localized basis sets) prohibitive for large systems. To address this computational cost, sparsity is exploited, and on-the-fly computation of integrals is employed for non-zero matrix elements. When these strategies are used, favourable scaling of mean-field-like methods makes them well-suited for large systems and molecular dynamics simulations.
Post-Hartree–Fock methods, such as configuration interaction [38], coupled cluster and multi-configurational approaches [39], achieve higher accuracy but at substantially higher computational costs. These methods typically involve a four-index transformation of integrals to the molecular orbital (MO) basis, which scales as but can be reduced to using low-rank approximations such as density fitting [40] or Cholesky decomposition [41]. The wave function is expressed in a basis of Slater determinants or configuration state functions (CSFs), whose size grows from to . Computational bottlenecks include finding the lowest roots of the Hamiltonian eigenvalue problem. The Hamiltonian matrix is generally too large to fit in memory, so iterative methods are used, such as Davidson’s algorithm [42], which is bound by sparse matrix–vector multiplications, with matrix elements evaluated on-the-fly using integrals expressed in the MO basis. Coupled cluster methods involve solving nonlinear equations [43], where tensor contractions dominate the computational cost [34]. Perturbative methods, such as Møller–Plesset perturbation theory (MP2), are well-suited for massive parallelism due to their non-iterative nature. Combined with the use of low-rank approximations, they can be used to study large systems with a high level of accuracy [44].
Kohn–Sham DFT and time-dependent DFT [45, 46] are widely used approaches for electronic structure calculations in the ground and excited state, though their accuracy depends on the choice of the approximate exchange–correlation functional and on the nature of the excitation (single or double excitation, Rydberg state, or charge transfer, for example). Green’s function approaches such as the GW and Bethe–Salpeter equation (BSE) methods are more expensive but can often provide more accurate band gaps than Kohn–Sham DFT and more accurate excitation energies than time-dependent DFT. However, doubly excited states and strongly correlated systems, for example, still require the use of wave function methods.
Quantum chemistry codes like GAMESS [47], NWChem [48], and Dalton [49] typically employ localized basis functions centred at the nuclear positions (Gaussian-type orbitals). These basis sets have the advantage of a compact representation. However, the computation of two-electron integrals often constitutes a substantial computational bottleneck. Developing efficient algorithms for integral evaluation remains an active area of research [50]. Although Gaussian basis functions are traditionally used for molecular simulations, they can also be applied to periodic systems, as implemented in PySCF [51, 52], CRYSTAL [53], and CRYSCOR [54], including in post-HF methodologies. Beyond this, versatile simulations of both molecular and condensed-matter systems are enabled by hybrid approaches that combine Gaussian and plane-wave basis sets, as in CP2K [55], or by using numerical atom-centered orbitals, as in FHI-aims [56].
For periodic systems, plane-wave-based codes such as VASP [57], Quantum Espresso [58], ABINIT [59] and CASTEP [60] are often used. Plane waves naturally accommodate periodic boundary conditions, making them ideal for studying bulk materials, surfaces and interfaces. Their computational bottleneck lies in the reliance on fast Fourier transforms (FFTs) to switch between real and reciprocal space. Communication overhead in large-scale parallel FFTs poses challenges for exascale architectures. Real-space grid-based codes such as Octopus [61], GPAW [62], SPARC [63], RMG [64], DFT-FE [65, 66], and PARSEC [67], provide an alternative approach by solving equations on a discrete grid. This is particularly advantageous for non-periodic and localized systems, allowing adaptive mesh refinement to enhance accuracy in critical regions while reducing computational costs. Despite their flexibility, these codes face their own challenges, primarily associated with efficient sparse matrix operations. Applying the sparse Hamiltonian matrix to the wave function demands highly optimized sparse matrix–vector multiplications and, for a very large scalability, sparse matrix algebra is a limiting factor before communication.
In periodic systems, wave function methods necessitate the use of a supercell to accurately account for electron–electron interactions. This approach significantly increases the size of the basis set, leading to substantial computational demands. Consequently, high-accuracy wave function calculations are rarely applied to periodic systems. An exception is the CC4S code [68], which enables periodic coupled cluster calculations and represents an important advance in this area. In condensed systems, for the computation of electronic excited states, the GW and BSE methods are preferred [69, 70]. Notable examples include BerkeleyGW [71], Exciting [72], and Yambo [73]. These specialized codes are typically used in conjunction with periodic DFT codes, using information from a preliminary ground-state calculation.
Real-space quantum Monte Carlo (QMC) methods, such as diffusion Monte Carlo, employ stochastic sampling to solve the many-body Schrödinger equation [74, 75, 76, 77]. These methods are inherently well-suited for massively parallel simulations and can achieve accuracies comparable to coupled cluster approaches. QMC is applicable to both periodic systems and molecular systems, using either plane-wave or localized basis sets. Its computational complexity and memory requirements are generally lower than those of other highly accurate wave-function-based methods, making QMC an attractive choice for larger systems. However, QMC methods have a high computational prefactor arising from the need for extensive sampling and the evaluation of complex wave functions. Key computational bottlenecks include the evaluation of the wave function at electron positions and the optimization of the wave function. Unlike deterministic methods, these challenges cannot be reduced to a few large linear algebra operations. Instead, QMC calculations involve a sequence of diverse computational kernels that handle relatively small data sizes but require repeated execution on a massive scale. Because of their specialized computational requirements, QMC methods are not typically implemented in mainstream quantum chemistry codes. Instead, they are supported by dedicated software packages such as QMCPACK [78, 79], TurboRVB [80], CHAMP [81], QMC=Chem [82], QWalk [83] and CASINO [84], which are interfaced with other codes to generate the trial wave functions needed for QMC calculations. Interestingly, the emergence of neural network-based wave functions has further expanded the scope of QMC methods. By making use of neural networks to parameterize the wave functions, these approaches can achieve highly accurate solutions to the electronic Schrödinger equation [85, 86, 87, 88, 89]. This innovation highlights the synergy between machine learning and quantum mechanical simulations.
Tackling the inherent trade-offs between accuracy and efficiency has driven researchers to increasingly explore artificial intelligence (AI) and machine learning (ML) as complementary tools. For instance, machine learning potentials trained on ab initio data can predict potential energy surfaces and forces with quantum-level accuracy at a significantly reduced computational cost [90, 91, 92, 93, 94, 95, 96, 97, 98, 99]. These advances have enabled efficient materials discovery and design through high-throughput calculations, supported by public datasets [100] and advanced optimization techniques [101, 102].
[H1] The transition to exascale
@@
The theoretical maximum speed of a computer is given by the total number of cores multiplied by their nominal speed. However, the actual speeds achieved can be far lower and need to be measured using benchmarks built on real applications with realistic data sets. The progress in HPC is marked by milestones tracked by the Top500 list [103], which monitors peak computational power through a benchmark known as LINPACK [104]. It measures the performance of systems in solving a dense linear system, a task optimized for high computational throughput with minimal data movement and communication. Remarkably, LINPACK achieves around 70% of the theoretical maximum speed. We have witnessed steady progress in computational power, culminating in crossing the exaflop barrier some time in 2022.
@@
However, LINPACK’s numbers, though impressive, are not fully representative of real-world performance. A more challenging and realistic benchmark is the High-Performance Conjugate-Gradient (HPCG) benchmark, which evaluates the ability of a computer to solve sparse linear systems and is therefore representative of a different type of numerical applications, for instance, based on finite elements and finite difference schemes [105]. The performance in this test is starkly different, typically approximately 20 times lower than LINPACK, using less than 5% of the nominal speed [106]. This contrast underscores the difference between idealized benchmarks and practical workloads.
@@
Although LINPACK and HPCG represent reasonable upper and lower bounds of performance, neither reflects the complexity of actual applications. The Gordon Bell Prize, awarded for outstanding achievements in HPC applications, provides a more relevant measure [107]. These applications, often solving real scientific problems, have demonstrated performance nearing the exaflop threshold [108, 109, 110, 111]. Achieving this requires monumental efforts: entire codebases are rewritten by interdisciplinary teams of physicists, chemists, mathematicians and computer scientists. New algorithms are developed, and implementations are tailored to exploit the specific features of the target architectures.
In contrast, many legacy codes, which comprise the bulk of real-world applications, lag far behind. Many such codes were not designed with exascale capabilities in mind and thus fall short of the exaflop performance. The challenge ahead lies in bridging this gap, ensuring that applications needing exascale capabilities can harness the full potential of exascale systems.
[H1] The changing hardware landscape
@@
The transition from petaflops to exascale computing marked a significant departure from the consistent and predictable trajectory of high-performance computing. Whereas the climb to petaflops was characterized by steady increases in processor core count and clock speeds, exascale computing brings a fundamental shift in hardware architectures and software development strategies. One of the most significant changes is the increasing adoption of heterogeneous architectures. Unlike the homogeneous central processing unit (CPU)-based systems that dominated petaflops, exascale systems heavily rely on accelerators like GPUs, which offer massive parallelism for specific tasks but come with unique challenges. This heterogeneous landscape presents a major hurdle for software developers as traditional algorithms and software stacks designed for CPUs need significant modifications to utilize the capabilities of GPUs efficiently.
@@
Before exploring the hardware evolution, it is crucial to identify the primary application areas driving advancements in hardware, as they fundamentally determine the level of investment and the trajectory of hardware development. Scientific HPC, including QM applications, remains a niche market and does not significantly influence major hardware innovations. Instead, scientific computing and exaflop architectures often repurpose software and hardware technology developed for other, larger markets. Two key segments, that are increasingly shaping the practices of the scientific community, are cloud and accelerator computing, driven primarily by their adoption in conventional companies and the field of AI, respectively. The general-purpose laptop segment also indirectly impacts high-end processor computing. For instance, Apple, a major customer of the Taiwan Semiconductor Manufacturing Company (TSMC), exerts significant pressure on high-end chip production for both cloud computing and AI applications due to their high volume demand, as these components share the same fabrication lines [112].
@@
On the CPU front, changes have been incremental over the past decade. Major CPU manufacturers (AMD, Intel, ARM) have converged towards a similar generic architecture: multicore processors based on out-of-order and superscalar technology. An important development has been the increase in the number of cores per processor, now exceeding 100. Although other technologies have emerged, none have achieved widespread adoption. Wider vectors (up to 512 bits) have been embraced by x86 architectures (initially by Intel, followed by AMD), whereas ARM continues to use shorter vectors (128 or 256 bits) without major performance penalties. Notably, memory technology has evolved more significantly, with very large level 3 caches using 3D stacking and High Bandwidth Memory (HBM) offering an alternative to the standard double data rate synchronous dynamic random access memory (DDR-SDRAM). The performance gains from these new memory technologies depend heavily on data access characteristics of the application.
Processing unit | Nvidia H100 SXMa | AMD EPYC™9754b |
Cores | 16896 | 128 |
Clock speed | 1.6 GHz | 2.25 GHz |
Release date | September 2022 | June 2023 |
FP64 | 34 TFlop/s | 6.9 TFlop/s |
FP64 (tensor core) | 67 TFlop/s | – |
FP32 | 67 TFlop/s | 13.8 TFlop/s |
FP32 (tensor core) | 989 TFlop/s | – |
FP16 (tensor core) | 1979 TFlop/s | – |
FP8 (tensor core) | 3958 TFlop/s | – |
Memory | 80 GB | max 6000 GB |
Memory Bandwidth | 3350 GB/s (HBM3) | 460.8 GB/s (DDR5) |
Interconnection Nvlink | 900 GB/s | |
Interconnection PCIe | 128 GB/s | |
Quantum X800 Infiniband network | 100 GB/s | 100 GB/s |
@@
On the accelerator front, the primary advance has been the massive increase in core count, now reaching thousands — nearly two orders of magnitude more than CPUs. The number of cores and peak performance depend on the floating-point (FP) format; typically, double-precision (DP) cores are half as numerous as single-precision (SP) cores. The shift toward narrower FP formats has been driven by AI, which tolerates reduced precision (16-bit, 8-bit, and even 4-bit formats are in use). Modern GPUs achieve remarkable performance using smaller FP formats and Tensor Cores, which are optimized for dense matrix multiplications. To illustrate these differences, Table 1 compares the computational performance of a high-performance CPU and a GPU released within the same time frame. Peak performance reflects execution speed only when computation is the bottleneck, not memory transfers or inter-node communications, which are too slow to keep up with the execution units. With standard DP operations (FP64), GPU peak performance is about five times that of a high-end CPU. When Tensor Cores are used, this increases to a factor of ten. For SP (FP32) operations, both CPU and GPU performance doubles, but with Tensor Cores, the GPU reaches 989 TFlop/s, over 70 times that of the CPU. Reducing precision to 8 bits (FP8) adds another factor of four. Whereas smaller FP formats can sometimes be used for standard FP32 and FP64 tasks, Tensor Cores are mainly of benefit to computations reliant on dense matrix multiplication. Achieving GPU peak performance is harder than on CPUs, as it requires high levels of parallelism. Consequently, matrices on the GPU must be much larger [115], or numerous, requiring batching algorithms. Furthermore, GPUs have less memory than CPUs [116] and large workloads necessitate memory transfers from the host to the accelerator, which must be managed in parallel with computations to prevent communication bottlenecks. We further elaborate below on the implications of these CPU and GPU performance differences for quantum simulations.
@@
Finally, on the supercomputing systems front, the changes have been rather radical. Using the Top500 list as a reference, most of the major changes in the supercomputing domain can be readily monitored. First, the amount of parallelism has drastically increased: the top runner in Top500 had 200,000 cores in June 2010, and now it has 11 million cores (1 million CPU and 10 million GPU cores; a 55 increase). The overall system organization has also radically changed: in 2010, GPUs were used in fewer than 40 systems among the 500, now over half of the systems are equipped with GPUs. Today’s top systems exhibit a heterogeneous architecture, integrating both CPUs and GPUs, though the bulk of computational capability predominantly stems from the GPUs. Moreover, the total power consumption of these systems has increased by an order of magnitude, rising from 3 MW to 30 MW, albeit with a notably irregular progression [103]. This increase underscores the substantial enhancements in system design, especially in terms of energy efficiency. This situation might change with the upcoming development of efficient processors and accelerator technology based on open-standard RISC-V instruction set architecture [117].
[H1] Increased fragmentation in the software stack
@@
To address portability, OpenCL was introduced. It provides an open standard for CPUs, GPUs and field programmable gate arrays (FPGAs) across vendors. However, it has not gained wide adoption in HPC due to performance and usability issues. Although OpenCL offers portability, its abstraction often results in suboptimal performance compared to specialized frameworks like CUDA [118]. Its lower-level programming model is also cumbersome, requiring substantial manual effort for memory management and kernel optimization. Fragmented implementations across vendors further undermine OpenCL’s portability, as performance and features can vary, forcing developers to fine-tune code for each platform.
The lack of an efficient standardized programming model for GPUs has resulted in a number of issues for developers. Firstly, it creates vendor lock-in, where codes written for one platform cannot easily be migrated to another without substantial code refactoring. For example, adopting Nvidia programming models involves considerable trade-offs between performance and portability across different hardware environments [119]. Another issue stemming from the fragmentation is the lack of compiler support for all programming models on all platforms. For instance, (Table 2), whereas CUDA is well supported on Nvidia GPUs through the CUDA Toolkit, its support on AMD and Intel hardware is limited and relies on indirect means like translation layers or third-party libraries and utilities [120, 121, 122, 123]. This forces developers to rely on vendor-specific compilers, which often lack the universality and feature-richness of their CPU counterparts.
Tool | C++ | Fortran | ||||
Intel | AMD | Nvidia | Intel | AMD | Nvidia | |
[…]
@@
The lack of standardization in GPU programming has also led to a variety of community-driven, higher-level models that aim to abstract away vendor-specific details and provide a more portable programming experience. Examples of such models include Kokkos [125], RAJA [126], Alpaka [127], and, to a lesser extent, hipSYCL [128] which is limited to AMD and Nvidia. These abstraction models often use vendor-native infrastructure in the background, enabling developers to write code that can be deployed on multiple platforms without substantial code changes. However, the support and standardization of these higher-level models can vary greatly, and relying on community-driven efforts for critical software infrastructures can introduce uncertainties and complexities in the long run. Comparing the performance of various GPU programming models on the LUMI supercomputer [129] reveals the potential of community-driven models like Kokkos in achieving portable performance but also highlights the need for further development and community support to ensure their long-term viability.
@@
Another important aspect of the software stack is the role of low-level libraries for key computational kernels, such as linear algebra, FFTs, and communication routines. These libraries play a critical role in achieving high performance and efficiency on exascale systems. However, the use of different libraries across different platforms can also contribute to software fragmentation. For example, whereas Nvidia provides the highly optimized cuBLAS library for linear algebra operations on their GPUs, AMD offers the RocBLAS library. These libraries, despite providing similar functionality, often have different APIs and performance characteristics, further hindering code portability.